# Nellie: A Neuro-Symbolic Inference Engine for Grounded, Compositional, and Explainable Reasoning

### v4

@@ Abstract
Our goal is a modern approach to answering questions via systematic reasoning where answers are supported by human interpretable proof trees grounded in an NL corpus of authoritative facts. Such a system would help alleviate the challenges of interpretability and hallucination with modern LMs, and the lack of grounding of current explanation methods (e.g., Chain-of-Thought). This paper proposes a new take on Prolog-based inference engines, where we replace handcrafted rules with a combination of neural language modeling, guided generation, and semiparametric dense retrieval. Our implementation, Nellie, is the first system to demonstrate fully interpretable, end-to-end grounded QA as entailment tree proof search, going beyond earlier work explaining known-to-be-true facts from text. In experiments, Nellie outperforms a similar-sized state-of-the-art reasoner (Tafjord et al., 2022) while producing knowledge-grounded explanations. We also find Nellie can exploit both semi-structured and NL text corpora to guide reasoning. Together these suggest a new way to jointly reap the benefits of both modern neural methods and traditional symbolic reasoning.
@@ 1 Introduction
Large language models (LLMs) are impressive at question-answering (QA), but it remains challenging to understand how answers systematically follow from authoritative information. This general opacity is a growing impediment to widespread use of LLMs, e.g., in critical applications such as medicine, law, and hiring decisions, where interpretability and trust are paramount. While there has been rapid progress on having LLMs generate systematic explanations for their answers, e.g., Chain-of-Thought (Wei et al., 2022), Entailer (Tafjord et al., 2022), or Maieutic Prompting (Jung et al., 2022), such explanations are not grounded in external facts and may include hallucinations (Ji et al., 2022). Rather, what would be desirable - and what this work pursues - is a system that systematically reasons over authoritative text: to support answers with human interpretable proof trees grounded in the text, while not requiring translation to an entirely symbolic formalism.
Figure 1: Given a query, Nellie performs a neuro-symbolic backward chaining search for proof trees whose leaves are grounded in a corpus of facts. First, it generates candidate decomposition rules guided by both retrieved facts and templates. Then, it recursively tries to prove the rules’ conditions via entailment from the corpus or further decomposition.
Our approach is to revisit the behavior of expert systems (Jackson, 1986; Metaxiotis et al., 2002). Expert systems are appealing for their explainable behavior: decisions are made by constructing a well-formed symbolic proof from explicit, formally represented knowledge authored by a knowledge engineer in consultation with a domain expert. However, as expert systems are known to be both expensive and brittle (Musen and Van der Lei, 1988), the AI community has turned to neuro-symbolic mechanisms that use large language models (LLMs) to reason over natural language (NL). Reasoning over NL does not require a symbolic formalism, and allows the use of inferentially powerful LLMs, but also loses the systematic, verifiable proofs that expert systems produced. This motivates our pursuit of a new way to jointly reap the benefits of both modern neural methods and traditional reasoning.
@@
As illustrated in Figure 2, various eXplainable QA (XQA) methods accomplish 2 of these criteria, but not all 3. Some methods, like Chain-of-Thought, extract inference chains or logically structured graphs fully from a LLM without grounding in verified knowledge; others, like ExplanationLP (Thayaparan et al., 2021), compose graphs of grounded facts, but do not show logical direction; finally, LAMBADA (Kazemi et al., 2023) achieves both direction and grounding, but is limited to simple domains with provided NL Horn rules and small sets of facts (1-2 dozen). In contrast, we desire a system that handles a much larger corpus (10K statements) and doesn’t need handcrafted rules.
@@
To achieve all three desiderata, we reuse the general inference framework of an expert system, but replace handcrafted rules with a combination of neural language modeling, guided generation, and semiparametric dense retrieval. We demonstrate this in a system called Nellie, the Neuro-Symbolic Large LM Inference Engine. Nellie leverages LLMs as proposal functions in the search of proof trees showing how a conclusion follows via entailment from an external corpus of NL statements. The “symbols” that our neuro-symbolic engine reasons over are free-form NL sentences. Rather than require a knowledge engineer to carefully write hundreds of inference rules as in the classical setting, Nellie employs an LLM as a dynamic rule generator  (DRG; Kalyanpur et al., 2021) to generate (rather than retrieve) candidate rules that decompose a hypothesis into subqueries that, if themselves proved recursively, will prove that hypothesis via composition (Figure 1). In this way, Nellie can answer the question posed by the classical expert system: “Does this statement systematically follow from my knowledge, or not?”.
@@
Nellie is built upon a backward chaining symbolic theorem prover written in Prolog, implemented using a few meta-rules specifying how inference should proceed. These include checking for entailment of a hypothesis against a retrieved fact, or decomposing it into a conjunction of subqueries to recursively prove. To treat NL sentences as if they were symbols in a purely symbolic search algorithm, we use an NLI-based form of weak unification (Sessa, 2002; Weber et al., 2019) between a queried statement and a corpus fact. This leads to interpretable proofs similar to the compositional entailment trees of prior work (e.g., EntailmentBank (Dalvi et al., 2021)), while tackling the additional challenge of multiple-choice QA.
@@
Nellie is designed to search across hundreds of trees whose leaves come from one of two types of knowledge: (a) a corpus of semi-structured text statements, such as NL renditions of database entries, or (b) a corpus of free-form NL sentences. Many correct proofs might exist for a given hypothesis, but much fewer will be fully groundable in the provided corpus, making the search harder than for ungrounded alternatives like Entailer (Tafjord et al., 2022) or COMET-Dynagen (Bosselut et al., 2021). To improve grounding, we introduce two guiding methods to boost the likelihood of generating rules whose conditions match against the available text: (1) For applications where a semi-structured corpus is available, Nellie leverages this structure via templates to help steer rule generation towards syntax that is more likely to match corpus entries. (2) For both free-form and semi-structured corpora, Nellie retrieves and conditions on statements to help steer generation towards trees grounded in them. Ablation experiments show these together and individually improve Nellie’s reasoning.
@@
Nellie expands upon work explaining known-to-be-true facts from text (SCSEARCH; Bostrom et al., 2022),(IRGR; Ribeiro et al., 2022) to perform QA, including searching for and comparing trees for conflicting options (both true and false). Experiments find Nellie outperforms a similar-sized SOTA reasoner (Entailer) while producing compositional trees showing how all decisions (top-level and intermediate) are grounded in corpus facts. Our contributions are thus:
@@
    An implementation, Nellie, [^1] Code will be released at https://github.com/JHU-CLSP/NELLIE that outperforms a similar-sized SOTA reasoner (Entailer-3B) while producing grounded trees. To our knowledge, this is the first system to perform grounded xQA as NL entailment tree search.
@@ 2 Related Work
Theorem Proving over Language A long-standing approach to reasoning over NL is to project it into a symbolic form, such as for QA (Green et al., 1961; Zelle and Mooney, 1996) or entailment (Bos and Markert, 2005). Provided the translation from NL to symbolic representation is accurate, one can leverage fast and scalable solutions for discrete symbolic inference (Riazanov and Voronkov, 2002; Kautz et al., 1992; Kautz and Selman, 1999). However, reliably translating broad-domain NL into an adequately expressive formalism for reasoning is a challenge (Schubert, 2015).
@@
Recent work explores methods of handling NL without needing to parse it to another formalism. Variants of neural theorem provers (NTPs; Rocktäschel and Riedel, 2017) such as NLProlog (Weber et al., 2019) reconcile NL facts with symbolic reasoning by learning embeddings for both the facts and the symbols in a theorem prover, then using weak unification to backward chain. (Kalyanpur et al., 2021) inject neural reasoning into a Boxer/Prolog-based symbolic reasoner via a special LM-calling predicate. (Arabshahi et al., 2021b) combine handwritten symbolic rules with neural symbol embeddings in order to classify conversational intents. Other work explores using LMs to emulate stepwise (Tafjord et al., 2021; Kazemi et al., 2023) or end-to-end (Clark et al., 2021; Picco et al., 2021) logical reasoning over small rulebases converted to NL.
@@
Systematic Explanation Generation Recent works have used LMs to generate NL reasoning graphs in support of an answer. “Chain-of-Thought” prompting (Wei et al., 2022; Kojima et al., 2022), elicits free-form inference hops from the LM before it generates an answer. Other text graph generation methods connect model-generated statements via common sense relations (Bosselut et al., 2021; Arabshahi et al., 2021a) or for/against influence (Madaan et al., 2021; Jung et al., 2022), though these are not knowledge-grounded and don’t address entailment (see Figure 2).
@@
The EntailmentBank dataset (Dalvi et al., 2021) has driven research towards the construction of explanation trees, challenging models to produce stepwise entailment proofs of a statement using a set of provided support facts. This direction builds upon works on explainable reasoning that build graphs from KB-retrieved sets of support statements, but stop short of showing their role in logical entailment (Pan et al., 2019; Jansen and Ustalov, 2019; Yadav et al., 2019; Valentino et al., 2022). Components of our framework are related to concurrent approaches for entailment tree construction (Bostrom et al., 2022; Hong et al., 2022). (Ribeiro et al., 2022) also use iterative retrieval-conditioned generation, and (Yang et al., 2022; Tafjord et al., 2022) also use entailment classifiers to filter proof steps. Few tree generation approaches consider this harder scenario of multiple-choice QA, opting instead to focus on the reconstruction task from support facts. (Tafjord et al., 2022) do consider the harder task. They propose a backward chaining QA system that generates entailment trees (not grounded in human-verified facts) using a search driven by internal model belief of factuality, complementary to Nellie’s use of guided and retrieval-conditioned generation to obtain knowledge grounding.
@@ 3 Background
Neural Predicates While most declarative predicates have meaning only in the context of user-defined inference axioms, others can call external neural modules that produce values for their arguments or determine the truth value of the predicate. A popular implementation of this paradigm is DeepProbLog (Manhaeve et al., 2018). We use this feature for defining neural model-invoking predicates. In the above example, we might train a seq2seq to produce other names for a disease $Y$, turning $\textsc{OtherName}(Y^{+},X^{-})$  [^2] In Prolog syntax, ‘ $+$ ’ denotes inputs, ‘ -- ’ outputs. into a neural predicate. This mechanism opens up the capacity to introduce externally-defined object symbols, e.g. seq2seq-generated NL, into the engine’s vocabulary.
Weak Unification In classical backward chaining, a unification operator assigns equivalence to two logical atoms; this requires atoms are of the same arity and have no unequal ground literals in the same argument position. Issues arise when literals are NL sentences, which can be syntactically distinct but semantically equivalent. To handle this, (Weber et al., 2019) propose a weak unification operator, which allows for the unification of any same-arity atoms regardless of conflicting symbols. [^3] Introducing weak unification greatly increases the search runtime, as one might try to unify any two symbols in the vocabulary at every recursive step. This poses a substantial challenge when applying the query grounding algorithm to a search space over NL. They estimate a unification score as the aggregation of pairwise similarity scores using a similarity function $\theta(s_{1},s_{2})\in[0,1]$. The score of the full proof is taken as the minimum of scores across all steps. In this work, we apply a similar aggregation; we say that a query fact $s_{1}$ “weakly unifies” with provenance fact $s_{2}$ with unification score equal to the confidence of an NLI model taking $s_{2}$ as the premise and $s_{1}$ as the hypothesis.
@@ 4 Overview of Approach
Depicted in Figure 3, our proposed framework is comprised of: An external corpus of facts (some $\{f_{1},\dots f_{n}\}$); a module that converts a QA pair to a hypothesis; an off-the-shelf theorem prover; and a suite of meta-axioms that use neural fact retrieval and dynamic rule generation modules to propose, verify, and score inferences. In our experiments, we consider one implementation of this framework that uses the corpus WorldTree (Xie et al., 2020), a set of 9K NL science facts that can interchangeably be considered as rows in 81 $n$-ary relational tables.
Question Conversion Given a multiple-choice question, the system converts each candidate answer into a hypothesis $h$ using a Question to Declarative Sentence model (QA2D; Demszky et al., 2018) (See §C). It then searches for a proof of $h$ against its knowledge base For each alternative, we enumerate $p$ proofs using a time-capped backward chaining search, and then take as the system’s answer the candidate with the overall highest-scoring proof.

### v5

@@ Abstract
Our goal is to develop a modern approach to answering questions via systematic reasoning where answers are supported by human interpretable proof trees grounded in an NL corpus of facts. Such a system would help alleviate the challenges of interpretability and hallucination with modern LMs, and the lack of grounding of current explanation methods (e.g., Chain-of-Thought). This paper proposes a new take on Prolog-based inference engines, where we replace handcrafted rules with a combination of neural language modeling, guided generation, and semiparametric dense retrieval. Our implementation, Nellie, is the first system to demonstrate fully interpretable, end-to-end grounded QA as entailment tree proof search, going beyond earlier work explaining known-to-be-true facts from text. In experiments, Nellie outperforms a similar-sized state-of-the-art reasoner while producing knowledge-grounded explanations. We also find Nellie can exploit both semi-structured and NL text corpora to guide reasoning. Together these suggest a new way to jointly reap the benefits of both modern neural methods and traditional symbolic reasoning.
@@ 1 Introduction
Large language models (LLMs) are impressive at question-answering (QA), but it remains challenging to understand how answers systematically follow from authoritative information. This general opacity is a growing impediment to widespread use of LLMs, e.g., in critical applications such as medicine, law, and hiring decisions, where interpretability and trust are paramount. While there has been rapid progress in having LLMs generate systematic explanations for their answers, e.g., Chain-of-Thought Wei et al. (2022), Entailer Tafjord et al. (2022), or Maieutic Prompting Jung et al. (2022), such explanations are not grounded in external facts and may include hallucinations Ji et al. (2022). Rather, what would be desirable - and what this work pursues - is a system that systematically reasons over authoritative text: to support answers with human interpretable proof trees grounded in the text, while not requiring translation to an entirely symbolic formalism.
Figure 1: Given a query, Nellie performs a neuro-symbolic backward chaining search for proof trees whose leaves are grounded in a corpus of facts. It generates candidate decomposition rules guided by retrieved facts or templates. Then, it recursively tries to prove rule conditions via entailment from the corpus or further decomposition.
Our approach is to revisit the behavior of expert systems Jackson (1986); Metaxiotis et al. (2002). Expert systems are appealing for their explainable behavior: decisions are made by constructing a well-formed symbolic proof from explicit, formally represented knowledge authored by a knowledge engineer in consultation with a domain expert. However, as expert systems are known to be both expensive and brittle Musen and Van der Lei (1988), the AI community has turned to neuro-symbolic mechanisms that use large language models to reason over natural language (NL). Reasoning over NL does not require a symbolic formalism and allows the use of inferentially powerful LLMs, but also loses the systematic, verifiable proofs that expert systems produced. This motivates our pursuit of a new way to jointly reap the benefits of both modern neural methods and traditional reasoning.
@@
As illustrated in Figure 2, various eXplainable QA (XQA) methods accomplish 2 of these criteria, but not all 3. Some methods, like Chain-of-Thought, generate inference chains or logically structured graphs from an LLM without grounding in verified knowledge. Others, like ExplanationLP Thayaparan et al. (2021), compose graphs of grounded facts, but do not show logical direction. LAMBADA Kazemi et al. (2023) achieves both direction and grounding but is limited to simple domains with provided NL Horn rules and sets of 1-2 dozen facts. In contrast, we desire a system that handles a larger corpus (10K statements) and doesn’t need handcrafted rules.
@@
To achieve all three desiderata, we reuse the general inference framework of an expert system, but replace handcrafted rules with a combination of neural language modeling, guided generation, and semiparametric dense retrieval. We demonstrate this in a system called Nellie, the Neuro-Symbolic Large LM Inference Engine. Nellie leverages LLMs as proposal functions in the search of proof trees showing how a conclusion follows via entailment from an external corpus of NL statements. The “symbols” that our neuro-symbolic engine reasons over are free-form NL sentences. Rather than require a knowledge engineer to carefully write hundreds of inference rules as in the classical setting, Nellie employs an LLM as a dynamic rule generator [DRG; Kalyanpur et al. (2021)] to generate (rather than retrieve) candidate rules that decompose a hypothesis into subqueries that, if themselves proved recursively, will prove that hypothesis via composition (Figure 1). In this way, Nellie can answer the question posed by the classical expert system: “Does this statement systematically follow from my knowledge, or not?”.
@@
Nellie is built upon a backward chaining symbolic theorem prover written in Prolog, implemented using a few meta-rules specifying how inference should proceed. These include checking for entailment of a hypothesis against a retrieved fact or decomposing it into a conjunction of subqueries to recursively prove. To treat NL sentences as if they were symbols in a purely symbolic search algorithm, we use a natural language inference-based form of weak unification Sessa (2002); Weber et al. (2019) between the hypothesis and a corpus fact. This produces interpretable proofs similar to the compositional entailment trees of prior work (e.g., EntailmentBank Dalvi et al. (2021)) while tackling the additional challenge of QA.
@@
Nellie is designed to search across hundreds of trees whose leaves come from one of two types of knowledge: (a) a corpus of semi-structured text statements, such as NL renditions of database entries, or (b) a corpus of free-form NL sentences. Many correct proofs might exist for a given hypothesis, but much fewer will be fully groundable in the provided corpus, making the search harder than for ungrounded alternatives like Entailer Tafjord et al. (2022) or COMET-Dynagen Bosselut et al. (2021). To improve grounding, we introduce two guiding methods to boost the likelihood of generating rules whose conditions match against the available text: (1) For applications where a semi-structured corpus is available, Nellie leverages this structure via templates to help steer rule generation towards syntax that is more likely to match corpus entries. (2) For both free-form and semi-structured corpora, Nellie retrieves and conditions on statements to help steer generation towards trees grounded in them. Ablation experiments show these together and individually improve Nellie’s reasoning.
@@
Nellie expands upon methods that ground known-to-be-true hypotheses to provided facts [SCSEARCH; Bostrom et al. (2022)], extending the paradigm to perform QA. This involves searching for and comparing trees for conflicting answer options. Experiments find Nellie outperforms a similar-sized state-of-the-art reasoner, Entailer, while producing compositional trees showing how decisions are grounded in corpus facts. Our contributions are thus:
@@
    An implementation, Nellie, [^1] Code found at https://github.com/JHU-CLSP/NELLIE. that outperforms a similar-sized SOTA reasoner (Entailer-3B) while producing grounded trees. To our knowledge, this is the first system to perform grounded XQA as NL entailment tree search.
@@ 2 Related Work
A long-standing approach to reasoning over NL is to project it into a symbolic form, such as for QA Green et al. (1961); Zelle and Mooney (1996) or entailment Bos and Markert (2005). Provided the translation from NL to symbolic representation is accurate, one can leverage fast and scalable solutions for discrete symbolic inference Riazanov and Voronkov (2002); Kautz et al. (1992); Kautz and Selman (1999). However, reliably translating broad-domain NL into an adequately expressive formalism for reasoning is a challenge Schubert (2015), though some have found success using LLMs to perform this semantic parsing task Wong et al. (2023); Lyu et al. (2023); Olausson et al. (2023); Pan et al. (2023); Ye et al. (2023).
@@
Recent work explores methods of handling NL without parsing it to another formalism. Some use LMs to generate proof steps in mathematical theorem proving Polu and Sutskever (2020); Welleck et al. (2022). Variants of neural theorem provers [NTPs; Rocktäschel and Riedel (2017)] such as NLProlog Weber et al. (2019) reconcile NL facts with symbolic reasoning by learning embeddings for the facts and symbols in a theory, then using weak unification to backward chain. Kalyanpur et al. (2021) inject neural reasoning into a Boxer/Prolog-based symbolic reasoner via a special LM-calling predicate. Arabshahi et al. (2021b) combine handwritten symbolic rules with neural symbol embeddings to classify conversational intents. Other work explores using LMs to emulate stepwise Tafjord et al. (2021); Kazemi et al. (2023) or end-to-end Clark et al. (2021); Picco et al. (2021) logical reasoning over small rulebases converted to NL. These approaches require both user-provided if/then rules to operate, while Nellie requires only facts and is thus applicable to domains in which rules are not available.
Modular Reasoning over NL
Nellie’s systematic reasoning is related to approaches that decompose problems into sequences of modular operations. Gupta et al. (2020) introduce a neural module network [NMN; Andreas et al. (2016)] for QA with modules for span extraction and arithmetic operations. Khot et al. (2021) introduce another NMN that decomposes questions into simpler ones answerable by existing models. These are part of a broader class of work decomposing multi-step reasoning using reasoning modules Khot et al. (2022); Saha et al. (2023a, b).
Systematic Explanation Generation
@@
Recent works have used LMs to generate NL reasoning graphs in support of an answer. “Chain-of-Thought” prompting Wei et al. (2022); Kojima et al. (2022), elicits free-form inference hops from the LM before it generates an answer. Other text graph generation methods connect model-generated statements via common sense relations Bosselut et al. (2021); Arabshahi et al. (2021a) or for/against influence Madaan et al. (2021); Jung et al. (2022), though these are not knowledge-grounded and don’t address entailment (see Figure 2).
@@
The EntailmentBank dataset Dalvi et al. (2021) has driven research towards the construction of explanation trees, challenging models to produce stepwise entailment proofs of a statement using a set of provided support facts. This direction builds upon works on explainable reasoning that build graphs from KB-retrieved sets of support statements, but stop short of showing their role in logical entailment Pan et al. (2019); Jansen and Ustalov (2019); Yadav et al. (2019); Valentino et al. (2022). Components of our framework are related to concurrent approaches for entailment tree construction Bostrom et al. (2022); Hong et al. (2022); Sprague et al. (2022). Ribeiro et al. (2022) also use iterative retrieval-conditioned generation, and Yang et al. (2022) also use entailment classifiers to filter proof steps. None of the above tree generation approaches consider this harder scenario of multiple-choice QA, opting instead to focus on the reconstruction task from support facts. Tafjord et al. (2022) do consider the harder task. They propose a backward chaining QA system, Entailer, that generates entailment trees (not grounded in human-verified facts) using a search grounded to the model’s internal beliefs, complementary to Nellie’s use of guided and retrieval-conditioned generation to obtain knowledge grounding. As with Nellie5.2), Entailer benefits from humans adding more useful statements to its available knowledge Dalvi Mishra et al. (2022). The difference between systems is highlighted by the number of generations (i.e. search nodes) considered by the two algorithms: while Entailer considers at most a couple dozen hypothesis decompositions, Nellie must consider hundreds or thousands to find one that is fully grounded.
Faithful Complex LLM Reasoning
Of the growing body of work on LLM-based “Chain-of-X” reasoning methods Xia et al. (2024), a portion considers ways to improve faithfulness to underlying knowledge and reduce LLM hallucinations (see Lyu et al. (2024) for an overview of faithful explanation methods). Nellie contributes to this space by providing a faithful-by-design algorithm that reasons based on logical hypothesis grounding. Other recent methods include Rethinking with Retrieval He et al. (2022), which reranks reasoning chains using a faithfulness score based on retrieved knowledge, and the forward-chaining Faithful Reasoning algorithm Creswell and Shanahan (2022), which follows “a beam search over reasoning traces” to answer a question by iteratively adding deductive inferences to a context of starting facts. Theirs is a different style search to our backward-chaining Nellie, not relying upon NLI for verifying logical connectedness and not scaling to larger knowledge bases.
@@ 3 Background
Neural Predicates
While most declarative predicates have meaning only in the context of user-defined inference axioms, others can call external neural modules that produce values for their arguments or determine the truth value of the predicate. A popular implementation of this paradigm is DeepProbLog Manhaeve et al. (2018), which we use to define LM-invoking predicates. In the above example, we might train a sequence-to-sequence (seq2seq) model to produce other names for a disease $Y$, turning $\textsc{OtherName}(Y^{+},X^{-})$  [^2] In Prolog syntax, ‘ $+$ ’ denotes inputs, ‘ -- ’ outputs. into a neural predicate. This mechanism creates the ability to introduce externally defined object symbols, e.g. seq2seq-generated NL, into the engine’s vocabulary.
Weak Unification
In classical backward chaining, a unification operator assigns equivalence to two logical atoms; this requires atoms to have the same arity and have no unequal ground literals in the same argument position. Issues arise when literals are NL sentences, which can be syntactically distinct but semantically equivalent. To handle this, Weber et al. (2019) propose a weak unification operator, which allows for the unification of any same-arity atoms regardless of conflicting symbols. [^3] Introducing weak unification greatly increases the search runtime, as one might try to unify any two symbols in the vocabulary at every recursive step. This poses a substantial challenge when applying the query grounding algorithm to a search space over NL. They estimate a unification score as the aggregation of pairwise similarity scores using a similarity function $\theta(s_{1},s_{2})\in[0,1]$. The score of the full proof is taken as the minimum of scores across all steps. In this work, we apply a similar aggregation; we say that a query fact $s_{1}$ “weakly unifies” with provenance fact $s_{2}$ with a unification score equal to the confidence of an NLI model taking $s_{2}$ as the premise and $s_{1}$ as the hypothesis.
@@ 4 Overview of Approach
Depicted in Figure 3, our framework is comprised of: an external corpus of facts (some $\{f_{1},\dots f_{n}\}$); a module that converts a QA pair to a hypothesis; an off-the-shelf theorem prover; and a suite of meta-axioms that use neural fact retrieval and dynamic rule generation modules to propose, verify, and score inferences. In our experiments, we consider one implementation of this framework that uses the corpus WorldTree Xie et al. (2020), a set of 9K NL science facts that can interchangeably be considered as rows in 81 $n$-ary relational tables.
Question Conversion
Given a multiple-choice question, the system converts each candidate answer into a hypothesis $h$ using a Question to Declarative Sentence model [QA2D; Demszky et al. (2018)] (See §C). It then searches for a proof of $h$ against its knowledge base. For each alternative, we enumerate $p$ proofs using a time-capped backward chaining search and then take as the system’s answer the candidate with the overall highest-scoring proof.